# Around the codifferential of products of p-forms and generalized interior products

### v2

@@ III A generalized interior product and its interplay with the Schouten-Nijenhuis bracket
and can be recovered using the graded Jacobi identity Eq. (10) and Eq. (14). In a similar vein setting $\alpha=\phi\in\Omega^{0}$ a function in Eq. (50) simplifies it to
@@
and can be recovered, then, directly from Eq. (48) and Eq. (14).
@@
in which the Schouten-Nijenhuis bracket appears naturally in the right-hand side. Equation (60) is obtained from Eq. (50) by specializing $\lambda=\flat v\in\Omega^{1}$ a $1$-form thus leading to
@@
$\llbracket\phi_{a},\cdots\llbracket\phi_{1},\alpha\rrbracket\cdots\rrbracket=(-1)^{a}i^{d\phi_{a}}\cdots i^{d\phi_{1}}\alpha=(-1)^{a}i^{d\phi_{1}\wedge\cdots\wedge d\phi_{a}}\alpha=(-1)^{a}(d\phi_{1}\wedge\cdots d\phi_{a},\alpha),$ (65)
@@
Finally, it can be interesting to use Tulczyjew[7], or Michor[8], differential operator $\Theta^{\alpha}=[i^{\alpha},d]_{a}$, cf. Eq. (58), fulfilling
$\Theta^{\phi}\alpha=-(d\phi)\wedge\alpha,$ $\phi\in\Omega^{0},\alpha\in\Omega^{a},$ (67)
$\Theta^{\lambda}\alpha=\mathscr{L}_{\sharp\lambda}\alpha,$ $\lambda\in\Omega^{1},\alpha\in\Omega^{a}$ (68)
$\Theta^{\alpha\wedge\beta}=i^{\beta}\Theta^{\alpha}+(-1)^{a}\Theta^{\beta}i^{\alpha}=(-1)^{ab}[i^{\alpha}\Theta^{\beta}+(-1)^{b}\Theta^{\alpha}i^{\beta}],$ $\alpha\in\Omega^{a},\beta\in\Omega^{b},$ (69)
@@
$i^{\llbracket\alpha,\beta\rrbracket}\gamma=(-1)^{a}i^{\beta}\Theta^{\alpha}\gamma+(-1)^{(a-1)(b-1)}\Theta^{\alpha}i^{\beta}\gamma,$ $\alpha\in\Omega^{a},\beta\in\Omega^{b},\gamma\in\Omega^{c},$ (71)
$\Theta^{\llbracket\alpha,\beta\rrbracket}=(-1)^{(a-1)(b-1)}\Theta^{\alpha}\Theta^{\beta}-\Theta^{\beta}\Theta^{\alpha},$ $\alpha\in\Omega^{a},\beta\in\Omega^{b}.$ (72)
@@ Appendix A Overview of needed formulas of differential geometry
Let $(M,g)$ be a smooth pseudo-riemannian manifold of dimension $n$, $TM$ and $T^{*}M$ being its tangent and cotangent fiber bundles, respectively. For $p$ an integer such that $1\leqslant p\leqslant n$ then $\Omega^{p}(M)\equiv\Omega^{p}=\bigwedge^{p}T^{*}M$ is the space of $p$-forms, smooth functions on $M$ are identified with zero-forms: $\Omega^{0}(M)=C^{\infty}(M,\mathbb{R})$.
The natural pairing between a $1$-form $\lambda$ and a vector $v$ is written $\langle\lambda,v\rangle$, the musical applications $\flat$ and $\sharp$ relate this pairing to the metric as $\langle\lambda,v\rangle=g(\sharp\lambda,v)$ and $g(u,v)=\langle\flat u,v\rangle.$ Of prime importance here are the ‘creator operator’ $j_{v}$, such that $j_{v}\alpha=(\flat v)\wedge\alpha$ with $v\in TM$ and $\alpha\in\Omega^{a}$, and the interior product $i_{v}$ (or insertion operator $i_{v}$) fulfilling an antisymmetry relation $i_{u}i_{v}=-i_{v}i_{u}$, a Leibniz identity $i_{v}\alpha\wedge\beta=(i_{v}\alpha)\wedge\beta+(-1)^{a}\alpha\wedge(i_{v}\beta)$ and interacting with $j_{v}$ as
@@
With the Hodge operator *, fulfilling $*1=\omega$ and $*\omega=\mathrm{sgn}(g)$ with $\omega$ the volume form, the product between $p$-forms is defined as
@@
see say App. A of Ref. 1 for additional identities. The generalized interior product stems from Eq. (75).
Appendix B Sketches of a few long computations
B.1 Eq. (50)
@@ Appendix B Sketches of a few long computations
B.2 Eqs. (54)–(61)
First notice that among the set of Eqs. (54)–(61) a few are known formulas rewritten as commutators, namely: Eq. (54) is Leibniz’ identity for the exterior derivative, Eq. (55) is Eq. (19), Eq. (56) is Leibniz’ identity for the Lie derivative, Eq. (57) is Eq. (21) which is obtained from Eq. (54) and (55). The remaining formulas, Eqs. (58)–(61), are found from the previous by conjugation with the Hodge operator: $x\beta\to*^{-1}x*\beta$.

### v3

@@ III.2 Combining the generalized product and the SN bracket
and can be recovered using the graded Jacobi identity Eqs. (11) and (15). In a similar vein setting $\alpha=\phi\in\Omega^{0}$ a function in Eq. (53) simplifies it to
@@
and can be recovered, then, directly from Eqs. (51) and (15).
@@
in which the Schouten–Nijenhuis bracket appears naturally in the right-hand side. Equation (63) is obtained from Eq. (53) by specializing $\lambda=\flat v\in\Omega^{1}$ a $1$-form thus leading to
@@
$\llbracket\phi_{a},\cdots\llbracket\phi_{1},\alpha\rrbracket\cdots\rrbracket=(-1)^{a}i^{d\phi_{a}}\cdots i^{d\phi_{1}}\alpha=(-1)^{a}i^{d\phi_{1}\wedge\cdots\wedge d\phi_{a}}\alpha=(-1)^{a}(d\phi_{1}\wedge\cdots\wedge d\phi_{a},\alpha),$ (68)
@@
Finally, it can be interesting to introduce Tulczyjew[7], or Michor[8], differential operator
$\Theta^{\alpha}=[i^{\alpha},d]_{a}=(i^{\alpha}d-(-1)^{a}di^{\alpha}),$ (70)
cf. Eq. (61) relating $\Theta$ to the Schouten–Nijenhuis bracket, fulfilling
$\Theta^{\phi}\alpha=-(d\phi)\wedge\alpha=-j^{d\phi}\alpha,$ $\phi\in\Omega^{0},\ \alpha\in\Omega^{a},$ (71)
$\Theta^{\lambda}\alpha=\mathscr{L}_{\sharp\lambda}\alpha,$ $\lambda\in\Omega^{1},\ \alpha\in\Omega^{a},$ (72)
$\Theta^{\alpha\wedge\beta}=i^{\beta}\Theta^{\alpha}+(-1)^{a}\Theta^{\beta}i^{\alpha}=(-1)^{ab}[i^{\alpha}\Theta^{\beta}+(-1)^{b}\Theta^{\alpha}i^{\beta}]$ (73)
[…]
@@
$i^{\llbracket\alpha,\beta\rrbracket}\gamma=(-1)^{a}i^{\beta}\Theta^{\alpha}\gamma+(-1)^{(a-1)(b-1)}\Theta^{\alpha}i^{\beta}\gamma$ (76)
[…]
$\Theta^{\llbracket\alpha,\beta\rrbracket}=(-1)^{(a-1)(b-1)}\Theta^{\alpha}\Theta^{\beta}-\Theta^{\beta}\Theta^{\alpha}=-[\Theta^{\beta},\Theta^{\alpha}]_{(a-1)(b-1)},$ $\alpha\in\Omega^{a},\ \beta\in\Omega^{b}.$ (78)
Out of these identities one can notice the special cases:
$\Theta^{\alpha\wedge\beta}\gamma=\begin{cases}\ [i^{\beta},\Theta^{\alpha}]\gamma=-i^{\llbracket\alpha,\beta\rrbracket}\gamma,\quad&\text{$a$ odd},\\[5.0pt] \ [i^{\beta},\Theta^{\alpha}]_{+}\gamma=i^{\llbracket\alpha,\beta\rrbracket}\gamma,&\text{$a$ even, $b$ odd},\\[5.0pt] \ i^{\llbracket\alpha,\beta\rrbracket}\gamma+2\Theta^{\alpha}i^{\beta}\gamma,&\text{$a$ even, $b$ even.}\end{cases}$ (79)
It also enables to recast Eq. (66) as
$[\mathscr{L}_{v},\delta]\beta=[\square,i_{v}]\beta-\Theta^{d\flat v}\beta,$ (80)
and in the special case of $v=\sharp\lambda$ a Killing vector it simplifies to
$[\square,i^{\lambda}]\beta=\Theta^{d\lambda}\beta.$ (81)
@@ Appendix A Overview of needed formulas of differential geometry
Let $(M,g)$ be a smooth pseudo-Riemannian manifold of dimension $n$, $TM$ and $T^{*}M$ being its tangent and cotangent fiber bundles, respectively. For $p$ an integer such that $1\leqslant p\leqslant n$ then $\Omega^{p}(M)\equiv\Omega^{p}=\smash{\bigwedge^{p}T^{*}M}$ is the space of $p$-forms, smooth functions on $M$ are identified with zero-forms: $\Omega^{0}(M)=C^{\infty}(M,\mathbb{R})$.
The natural pairing between a $1$-form $\lambda$ and a vector $v$ is written $\langle\lambda,v\rangle$, the musical applications $\flat$ and $\sharp$ relate this pairing to the metric as $\langle\lambda,v\rangle=g(\sharp\lambda,v)$ and $g(u,v)=\langle\flat u,v\rangle.$ Of prime importance here are the “creator operator” $j_{v}$, such that $j_{v}\alpha=(\flat v)\wedge\alpha$ with $v\in TM$ and $\alpha\in\Omega^{a}$, and the interior product $i_{v}$ (or insertion operator $i_{v}$) fulfilling an antisymmetry relation $i_{u}i_{v}=-i_{v}i_{u}$, a Leibniz identity
$i_{v}\alpha\wedge\beta=(i_{v}\alpha)\wedge\beta+(-1)^{a}\alpha\wedge(i_{v}\beta),\quad\alpha\in\Omega^{a},\ \beta\in\Omega^{b},$ (82)
and interacting with $j_{v}$ as
@@
On an oriented manifold with the Hodge operator *, fulfilling $*1=\omega$, $*\omega=\mathrm{sgn}(g)$ with $\omega$ the volume form, and $**\alpha=\mathrm{sgn}(g)(-1)^{a(n+1)}\alpha$ for $\alpha\in\Omega^{a}$, the product between $p$-forms is defined as
@@
see, say, App. A of Ref. 1 for additional identities. The generalized interior product stems from Eq. (85).
Appendix B Proofs and sketches of a few long computations
B.1 Proof of Eq. (9)
Proof.
First, notice that both $\llbracket\lambda,\beta\rrbracket-\mathscr{L}_{\sharp\lambda}\beta$ and $(\mathscr{L}_{\sharp\lambda}\tilde{g})^{mn}j_{m}i_{n}\beta$ vanish on functions (i.e. 00-forms) and agree on $1$-forms. Indeed, on a $1$-form $\tau\in\Omega^{1}$ one has $\llbracket\lambda,\tau\rrbracket-\mathscr{L}_{\sharp\lambda}\tau=\flat\mathscr{L}_{\sharp\lambda}\sharp\tau-\mathscr{L}_{\sharp\lambda}\tau$ but
$\flat\mathscr{L}_{\sharp\lambda}\sharp\tau$ $=\flat\mathscr{L}_{\sharp\lambda}C(\tilde{g},\tau)=\flat C(\mathscr{L}_{\sharp\lambda}\tilde{g},\tau)+\flat C(\tilde{g},\mathscr{L}_{\sharp\lambda}\tau)$
$=\flat C(\mathscr{L}_{\sharp\lambda}\tilde{g},\tau)+\mathscr{L}_{\sharp\lambda}\tau=(\mathscr{L}_{\sharp\lambda}\tilde{g})^{mn}j_{m}i_{n}\tau+\mathscr{L}_{\sharp\lambda}\tau,$ (87)
where $C$ is the contraction of a symmetric 2-vector with a 1-form.
Furthermore, we note that $\llbracket\lambda,\beta\rrbracket-\mathscr{L}_{\sharp\lambda}\beta$ and $(\mathscr{L}_{\sharp\lambda}\tilde{g})^{mn}j_{m}i_{n}\beta$ are both derivatives on $p$-forms fulfilling the same Leibniz identity (derivation of degree 0 on the graded algebra $\Omega$). Indeed, the same Leibniz identity holds for the three terms above: for $\llbracket\lambda,\cdot\rrbracket$ thanks to Eq. (6), for $\mathscr{L}_{\sharp\lambda}$ it is the usual one and for $j_{m}i_{n}$ thanks to Eq. (82) and to the fact that $j_{m}\alpha\wedge\beta=(-1)^{a}\alpha\wedge j_{m}\beta$ for $\alpha\in\Omega^{a}$ and $\beta\in\Omega^{b}$. Finally, these two derivations of degree zero are equal since they agree on 0-forms and 1-forms. ∎
B.2 Proof of Prop. 1
Proof.
Expanding first $\alpha$, using Eq. (13), brings about
using the skew-symmetry of the coefficient. Then, permuting the arguments in the brackets and using again this formula, gives rise to
B.3 Proof of Prop. 3
Proof.
First let $a=0$, that is $\alpha=\phi\in\Omega^{0}$ be a function and $\beta\in\Omega^{b}$ a $b$-form, then
which follows from the definition of the codifferential $\delta\beta=(-1)^{b}*^{-1}d*\beta$, and that, on the one hand, one has $\llbracket\phi,\beta\rrbracket=-i^{d\phi}\beta$ (cf. Eq. (15)) and, on the other hand, one has $i^{d\phi}\beta=\phi\delta\beta-\delta(\phi\beta)$, combining both equations proves the formula for $a=0$.
Now let $a=1$, that is $\alpha$ being a $1$-form, and $\beta\in\Omega^{b}$ a $b$-form, we then have
$\delta(\alpha\wedge\beta)=-\alpha\wedge\delta\beta-*^{-1}\mathscr{L}_{\sharp\alpha}*\beta,$ (92)
by commuting $j^{\alpha}$ with * in $\delta$ with Eq. (86) and using Cartan formula to recast $di^{\alpha}$ as $\mathscr{L}_{\sharp\alpha}-i^{\alpha}d$ and finally commuting the remaining $i^{\alpha}$ with $*^{-1}$ with Eq. (85). Then owing to Eq. (19) this reads as
$\delta(\alpha\wedge\beta)=-\alpha\wedge\delta\beta-(\llbracket\alpha,\beta\rrbracket-(\delta\alpha)\beta)=(\delta\alpha)\beta-\alpha\wedge\delta\beta-\llbracket\alpha,\beta\rrbracket,$ (93)
thus proving the formula for $a=1$.
The rest of the proof relies on an induction on $a$. Indeed, assume that the formula holds at an arbitrary $a$, which is already the case for $a=0$ and $a=1$ (cf. Eqs. (91) and (93)), and consider $\alpha\in\Omega^{a+1}$ a $(a+1)$-form. Let $\beta\in\Omega^{b}$ be a $b$-form, then
$\delta(\alpha\wedge\beta)$ $=\delta(\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge e^{m_{a+1}}\wedge\beta)$
$=(\delta\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}})\wedge e^{m_{a+1}}\wedge\beta$
$\qquad+(-1)^{a}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\delta(e^{m_{a+1}}\wedge\beta)$
$\qquad+(-1)^{a}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\wedge\beta\rrbracket$
$=(\delta\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}})\wedge e^{m_{a+1}}\wedge\beta$
$\qquad+(-1)^{a}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\delta(e^{m_{a+1}})\wedge\beta$
$\qquad+(-1)^{a+1}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge e^{m_{a+1}}\wedge\delta\beta$
$\qquad+(-1)^{a+1}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\llbracket e^{m_{a+1}},\beta\rrbracket$
$\qquad+(-1)^{a}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\wedge\beta\rrbracket$
$=(\delta\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge e^{m_{a+1}})\wedge\beta$
$\qquad+(-1)^{a+1}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\rrbracket\wedge\beta$
$\qquad+(-1)^{a+1}\alpha\wedge\delta\beta$
$\qquad+(-1)^{a+1}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\llbracket e^{m_{a+1}},\beta\rrbracket$
$\qquad+(-1)^{a}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\wedge\beta\rrbracket$
$=(\delta\alpha)\wedge\beta+(-1)^{a+1}\alpha\wedge\delta\beta$
$\qquad+(-1)^{a+1}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\rrbracket\wedge\beta$
$\qquad+(-1)^{a+1}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\llbracket e^{m_{a+1}},\beta\rrbracket$
$\qquad+(-1)^{a}\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},e^{m_{a+1}}\wedge\beta\rrbracket$
$=(\delta\alpha)\wedge\beta+(-1)^{a+1}\alpha\wedge\delta\beta$
$\qquad+(-1)^{a+1}\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}}\wedge\llbracket e^{m_{a+1}},\beta\rrbracket$
$\qquad+e^{m_{a+1}}\wedge\llbracket\alpha_{m_{1}\cdots m_{a+1}}e^{m_{1}}\wedge\cdots\wedge e^{m_{a}},\beta\rrbracket$
$=(\delta\alpha)\wedge\beta+(-1)^{a+1}\alpha\wedge\delta\beta+(-1)^{a+1}\llbracket\alpha,\beta\rrbracket,$
in which the combinatorial factor $1/(a+1)!$ has been incorporated in $\alpha_{m_{1}\cdots m_{a+1}}$ as it plays no role here, then the recurrence hypothesis is used in the first to second line, Eq. (93) is used on the second term in the second to third line, the recurrence hypothesis is again used to recognize $(\delta\alpha)\wedge\beta$ and thus creating a SN bracket in the third to fourth line. The fifth line is a rewriting of the fourth putting first the sought for terms and keeping the SN brackets to reshape as additional terms. Equation (6) is used on the last term of the fifth line which simplifies partially with the first SN bracket. Finally Eq. (13) is used to recognize in the last two terms $(-1)^{a+1}\llbracket\alpha,\beta\rrbracket$. ∎
B.4 Derivation of Eq. (53)
@@ Appendix B Proofs and sketches of a few long computations
B.5 Derivation of Eqs. (57)–(64)
First, notice that among the set of Eqs. (57)–(64) a few are known formulas rewritten as commutators, namely: Eq. (57) is Leibniz identity for the exterior derivative, Eq. (58) is Eq. (22), Eq. (59) is Leibniz identity for the Lie derivative, Eq. (60) is Eq. (29) which is obtained from Eqs. (57) and (58). The remaining formulas, Eqs. (61)–(64), are found from the previous by conjugation with the Hodge operator: $x\beta\to*^{-1}x*\beta$.
B.6 Derivation of Eqs. (71)–(78)
The first few equations in Eqs. (71)–(78) are straight consequences of the definition (70) of $\Theta^{\alpha}$ and that of the signed commutator along with Cartan and Leibniz identities, taking into account $d^{2}=0$ and $i^{\alpha\wedge\beta}=i^{\beta}i^{\alpha}=(-1)^{ab}i^{\alpha}i^{\beta}$ (cf. Eq. (44)). Equation (76) relies on Eq. (53) and on Eq. (44) to recast, say, $i^{\beta}d$ as $\Theta^{\beta}+(-1)^{b}di^{\beta}$ thus leading to the result and its various expressions. Establishing Eq. (78) starts with the definition (70) of $\Theta^{\alpha}$ and by using directly Eq. (76), then using Eq. (75) to commute $\Theta$ and $d$ yields the result.